You have a scriptbase text Li end text end scriptbase presubscript 3 presuperscript 7 atom (3 protons) which is doubly ionized such that it has just one remaining electron. (scriptbase text Li end text to the power of 2 plus end exponent end scriptbase presubscript 3 presuperscript 7)

How much energy, in electron volts, does it take to remove the remaining electron if it starts in the n = 1 orbital? (to 2 s.f and in eV)
1
Expert's answer
2020-06-01T13:12:44-0400

The ionization energy required is


"E=13.6\frac{Z^2}{n^2}=13.6\frac{3^2}{1^2}=122\text{ eV}."